OpenStudy (anonymous):

how do you solve 6x^2-8x-7=0

OpenStudy (anonymous):

use quadratic equation form u know how to solve it?

OpenStudy (anonymous):

( -b + - Square -b^2 - (4*a*c)): 2a

OpenStudy (anonymous):

a=6 b=-8 c=-7

OpenStudy (anonymous):

\[(-(-8)\pm \sqrt{-8^{2}-(4*6*(-7)})\div(2*6)\]\[(8\pm \sqrt{64+168})\div6\]\[(8\pm \sqrt{232})\div6\]\[x=(1/6)(4\pm58)\]